An Account of the Life of George Berkeley,

D.D. Late Bishop of Cloyne in Ireland

Joseph Stock

Published 1776

INTRODUCTION.

There seems to be an odd fatality attending upon some of the first characters in the republic of letters, that the very celebrity they had deservedly acquired amongst their contemporaries has prevented an accurate knowledge of their lives from descending to posterity. A writer, distinguished by uncommon abilities, more especially if that writer has acted a busy part on the stage of life is so frequently the subject of conversation, that for some years after his removal the memories of those who knew him are thought to be sufficiently secure repositories of his fame; till by degrees the fading materials on which his actions were written moulder away, and curiosity begins precisely at the point of time when the means of gratifying it are lost. How nearly this hath been the case of the excellent prelate whose life, character, and writings we have here attempted to describe, the reader will be able to form a judgment when he is assured, that in more than twenty years which have elapsed since the death of Bishop BERKELEY, no account of him hath yet been offered to the public that was not either void of truth, or extremely inaccurate and defective. Neither is this intended as a censure upon such as wrote from what information they could collect, and probably thought any account, however imperfect, of so extraordinary a person, better than none: it is only offered as an excuse for the present undertaking, to which the author is conscious he brings no other qualification than knowledge of the truth of every fact he relates, and an entire freedom from prejudice. Particular acquaintance with the family and friends of Bishop Berkeley has put him in possession of the first; the course itself of the following narrative will best shew, whether he has any just pretensions to the latter.

LIFE OF BISHOP BERKELEY.

DR. GEORGE BERKELEY, the learned and ingenious bishop of Cloyne in Ireland, was a native of that kingdom, and the son of William Berkeley of Thomastown, in the county of Kilkenny, whose father went over to Ireland after the Restoration (the family having suffered greatly for their loyalty to Charles I.) and there obtained the collectorship of Belfast.

Our author was born March 12, 1684, at Kilcrin near Thomastown, received the first part of his education at Kilkenny school under Dr. Hinton, and was admitted a pensioner of Trinity College, Dublin, at the age of fifteen, under the tuition of Dr. Hall. He was chosen fellow of that college June 9, 1707, having previously sustained with honour the very trying examination, which the candidates for that preferment are by the statutes required to undergo.

The first proof he gave of his literary abilities was (A) Arithmetica absque Algebra aut Euclide demonstrata, which, from the preface, he appears to have written before he was twenty years old, though he did not publish it till 1707. It is dedicated to Mr. Palliser, son to the archbishop of Cashel, and is followed by a Mathematical Miscellany, containing some very ingenious observations and theorems inscribed to his pupil Mr. Samuel Molyneux, a gentleman of whom we shall have occasion to make further mention presently, and whose father was the celebrated friend and correspondent of Mr. Locke.

His (B) Theory of Vision was published in 1709, and the (C) Principles of human Knowledge appeared the year after. The airy visions of romances, to the reading of which he was much addicted, disgust at the books of metaphysics then received in the University, and that inquisitive attention to the operations of the mind which about this time was excited by the writings of Mr. Locke and Father Malebranche, probably gave birth to his disbelief of the existence of matter.

In 1712, the principles inculcated in Mr. Locke’s Two Treatises of Government, seem to have turned his attention to the doctrine of passive obedience; in support of which he printed the substance of three Common-places delivered by him that year in the college chapel, a work which afterwards did him some injury in his fortune. For being presented by Mr. Molyneux abovementioned to their late majesties, then Prince and Princess of Wales (whose secretary Mr. Molyneux had been at Hanover) he was by them recommended to Lord Galway for some preferment in the church of Ireland. But Lord Galway having heard of those sermons, represented him as a Jacobite, an impression which Mr. Molyneux as soon as he was apprised of it, took care to remove from the minds of their highnesses by producing the work in question, and shewing that it contained nothing but principles of loyalty to the present happy establishment. This was the first occasion of our Author’s being known to Queen Caroline.

In February 1713 he crossed the water, and published in London a further defence of his celebrated system of immaterialism, in Three dialogues between Hylas and Philonous. Acuteness of parts and a beautiful imagination were so conspicuous in his writings, that his reputation was now established, and his company was courted, even where his opinions did not find admission. Two gentlemen of opposite principles concurred in introducing him to the acquaintance of the learned and the great, Sir Richard Steele, and Dr. Swift. He wrote several papers in the Guardian for the former, and at his house became acquainted with Mr. Pope, with whom he continued to live in strict friendship during his life. Dean Swift, besides Lord Berkeley of Stratton (D) (to whom our author dedicated his last published dialogues between Hylas and Philonous) and other valuable acquaintance, recommended him to the celebrated Earl of Peterborough, who being appointed Embassador to the King of Sicily and to the other Italian states, took Mr. Berkeley with him in quality of chaplain and secretary, in November 1713.

At Leghorn, his lordship’s well known activity induced him to disencumber himself of his chaplain and the greatest part of his retinue, whom he left in that town for upwards of three months, while he discharged the business of his embassy in Sicily, as our author informs his friend Pope in the conclusion of a complimentary letter^{1} addressed to that poet on the Rape of the Lock, dated Leghorn, May 1, 1714. It may not be amiss to record a little incident that befel Mr. Berkeley in this city, with the relation of which he used sometimes to make himself merry among his friends. Basil Kennett, the author of the Roman Antiquities, was then chaplain to the English factory at Leghorn, the only place in Italy where the English service is tolerated by the government, which favour had lately been obtained from the Grand Duke at the particular instance of Queen Anne. This gentleman requested Mr. Berkeley to preach for him one Sunday. The day following, as Berkeley was sitting in his chamber, a procession of priests in surplices, and with all other formalities, entered the room, and without taking the least notice of the wondering inhabitant, marched quite round it, muttering certain prayers. He fears immediately suggested to him that this could be no other than a visit from the Inquisition, who had heard of his officiating before heretics without licence, the day before. As soon as they were gone, he ventured with much caution to enquire into the cause of this extraordinary appearance, and was happy to be informed, that this was the season appointed by the Romish calendar for solemnly blessing the houses of all good catholics from rats and other vermin; a piece of intelligence which changed his terror into mirth.

1Pope’s Works.

He returned to England with Lord Peterborough in August 1714, (E) and his hopes of preferment through this channel expiring with the fall of Queen Anne’s ministry, he some time after embraced an advantageous offer made him by Dr. St. George Ashe, Bishop of Clogher, and late Provost of Trinity College, Dublin, of accompanying his son, Mr. Ashe (who was heir to a very considerable property) in a tour through Europe.

At Paris, having now more leisure than when he first passed through that city, Mr. Berkeley took care to pay his respects to his great rival in metaphysical sagacity, the illustrious Pere Malebranche. He found this ingenious father in his cell, cooking in a small pipkin a medicine for a disorder with which he was then troubled, an inflammation on the lungs. The conversation naturally turned on our author’s system of which the other had received some knowledge from a translation just published. But the issue of this debate proved tragical to poor Malebranche.—In the heat of disputation he raised his voice so high, and gave way so freely to the natural impetuosity of a man of parts and a Frenchman, that he brought on himself a violent increase of his disorder, which carried him off a few days after^{2}.

In this second excursion abroad Mr. Berkeley employed upwards of four years; and besides all those places which are usually visited by travellers in what is called the grand tour, his curiousity carried him to some that are less frequented. In particular he travelled over Apulia, the Tarentese (from which he wrote an accurate and entertaining account of the tarantula to Dr. Friend), Calabria, and the whole island of Sicily. This last country engaged his attention so strongly, that he had with great industry compiled very considerable materials for a natural history of the island: but by an unfortunate accident these, together with a journal of his transactions there, were lost in the passage to Naples, nor could he be prevailed upon afterwards to recollect and commit those curious particulars again to paper (F). What an injury the literary world has sustained by this mischance, may in part be collected from the specimen he has left of his talent for lively description, in his letter to Mr. Pope (G) concerning the island of Inarime (now Ischia, in the bay of Naples) dated Naples, October 22, 1717; and in another from the same city to Dr. Arbuthnot, giving an account of an eruption of Mount Vesuvius, which he had the good fortune to have more than one opportunity of examining very minutely.

2He died October 13, 1715. Dict. hist. portatif d’Advocat.

On his way homeward he drew up at Lyons a curious tract (H) De motu, which he sent to the royal academy of sciences at Paris, the subject being proposed by that assembly, and committed it to the press shortly after his arrival in London in 1721. But from these abstruse speculations he was drawn away for a while by the humanity of his temper and concern for the public welfare. It is well known what miseries the nation was plunged into by the fatal South Sea scheme in 1720. Mr. Berkeley felt for his country and British neighbours groaning under these calamitous distresses, and in that spirit employed his talents in writing An Essay towards preventing the Ruin of Great Britain, printed London 1721.

His travels had so far improved his natural politeness, and added such charms to his conversation, that he found a ready admission into the best company in London. Among the rest, Mr. Pope introduced him to Lord Burlington, who conceived a high esteem for him on account of his great taste and skill in architecture, an art of which his Lordship was an excellent judge and patron, and which Mr. Berkeley had made his particular study while in Italy. By this nobleman he was recommended to the Duke of Grafton, Lord Lieutenant of Ireland, who took him over to Ireland as one of his chaplains in 1721, after he had been absent from his native country more than six years. He had been elected a senior fellow of his college in July 1717, and now took the degrees of bachelor and doctor in divinity, November 14, 1721.

The year following his fortune received a considerable increase from a very unexpected event. On his first going to London in the year 1713, Dean Swift introduced him to the family of Mrs. Esther Vanhomrigh (the celebrated Vanessa), and took him often to dine at her house. Some years before her death, Vanessa had removed to Ireland, and fixed her residence at Cellbridge, a pleasant village in the neighbourhood of Dublin, most probably with a view of often enjoying the company of a man, for whom she seems to have entertained a very singular attachment. But finding herself totally disappointed in this expectation, and discovering the Dean’s connection with Stella, she was so enraged at his infidelity, that she altered her intention of making him her heir, and left the whole of her fortune, amounting to near 8000l. to be divided equally between two gentlemen whom she named her executors, Mr. Marshal, a lawyer, afterwards Mr. Justice Marshal, and Dr. Berkeley, S.F.T.C.D. The doctor received the news of this bequest from Mr. Marshal with great surprise, as he had never once seen the lady who had honoured him with such a proof of her esteem, from the time of his return to Ireland to her death. It was fortunate however for Swift, that Berkeley was chosen one of her executors: for, in consequence of this trust, the whole correspondence between Cadenus and Vanessa, as well as the poem that goes under that name, fell into the hands of a man who had so much tenderness for his friend’s reputation as to burn the letters immediately, though he saw nothing in the verses that should hinder their publication. From some fragments that have since got into print (probably hasty extracts taken by such as had seen them before they came into Dr. Berkeley’s hands) it appears, that if there was nothing criminal, there was at least a warmth in those letters, that justifies Dr. Berkeley’s delicacy in suppressing them. Mr. Marshal, the other executor, had entered so far into the resentment of his benefactress against Swift, and was besides so attached to that Bettesworth whom the Dean’s satiric muse has immortalized, that he was not without difficulty prevailed upon to give his consent to their suppression.

May 18, 1724, Dr. Berkeley resigned his fellowship, being promoted by his patron the Duke of Grafton to the deanery of Derry, worth 1100l. per annum. In the interval between this removal and his return from abroad, his mind had been employed in conceiving that benevolent project, which alone entitles him to as much honour as all his learned labours have procured him, the Scheme for converting the savage Americans to Christianity, by a College to be erected in the Summer islands, otherwise called the isles of Bermuda. He published a proposal for this purpose, London 1725, and offered to resign his own opulent preferment, and to dedicate the remainder of his life to the instructing the youth in America, on the moderate subsistence of 100l. yearly. Such was the force of this disinterested example, supported by the eloquence of an enthusiast for the good of mankind, that three junior fellows of Trinity College, Dublin, the Reverend William Thompson, Jonathan Rogers, and James King, masters of arts, consented to take their fortunes with the author of the project, and to exchange for a settlement in the Atlantic ocean at 40l. per annum, all their prospects at home; and that too at a time, when a fellowship of Dublin College was supposed to place the possessor in a fair point of view for attracting the notice of his superiors both in the church and state.

Dr. Berkeley however was not so ill acquainted with the world, as to rest the success of his application to the ministry entirely on the hope his scheme afforded of promoting national honour and the cause of Christianity: his arguments were drawn from the more alluring topic of present advantage to the government. Having with much industry acquired an accurate knowledge of the value of certain lands (K) in the island of St. Christopher’s, yielded by France to Great Britain at the treaty of Utrecht, which were then to be sold for the public use, he undertook to raise from them a much greater sum than was expected, and proposed that a part of the purchase money should be applied to the erecting of his college. He found means, by the assistance of a Venetian of distinction, the Abbé Gualtieri (or Altieri) with whom he had formed an acquaintance in Italy, to carry this proposal directly to King George I. (L) who laid his commands on Sir Robert Walpole to introduce and conduct it through the House of Commons. His Majesty was further pleased to grant a charter for erecting a college by the name of St. Paul’s College in Bermuda, to consist of a president and nine fellows, who were obliged to maintain and educate Indian scholars at the rate of 10l. per annum for each. The first president, Dr. George Berkeley, and first three fellows named in the charter (being the gentlemen above-mentioned) were licenced to hold their preferments in these kingdoms till the expiration of one year and a half after their arrival in Bermuda. The Commons, May 11, 1726, voted, “That an humble address be presented to his majesty, that out of the lands in St. Christopher’s, yielded by France to Great Britain by the treaty of Utrecht, his majesty would be graciously pleased to make such grant for the use of the president and fellows of the college of St. Paul, in Bermuda, as his majesty shall think proper.” The sum of 10,000l. was accordingly promised by the minister, and several private subscriptions were immediately raised for promoting “so pious an undertaking,” as it it stiled in the king’s answer^{3} to this address. Such a project of success in the favourite object of his heart, drew from our author a beautiful copy of verses (M) in which another age perhaps will acknowledge the old conjunction of the prophetic character with that of the poet to have again taken place.

In the mean time the Dean entered into a marriage, August 1, 1728, with Anne, the eldest daughter of the Right Honourable John Forster, Esq; speaker of the Irish house of Commons. This engagement however was so far from being any obstruction to his grand undertaking, that he actually set sail in the execution of it for Rhode Island about the middle of September following. He carried with him his lady, a Miss Handcock, two gentlemen of fortune, Mess. James and Dalton, a pretty large sum of money of his own property, and a collection of books for the use of his intended library (N). He directed his course to Rhode Island, which lay nearest to Bermuda, with a view to purchasing lands on the adjoining continent as estates for the support of his college; having a positive promise from those in power, that the parliamentary grant should be paid him as soon as ever such lands should be pitched upon and agreed for. The Dean took up his residence at Newport in Rhode Island, where his presence was a great relief to a clergyman of the church of England established in those parts, as he preached every Sunday, and was indefatigable in pastoral labours during the whole of his stay there, which was near two years.

When estates had been agreed for, it was fully expected that the public money would, according to grant, be immediately paid as the purchase of them. But the minister had never heartily embraced the project, and parliamentary influence had by this time interposed, in order to divert the grant into another channel. The sale of the lands in St. Christopher’s, it was found, would produce 90,000l. Of this sum 80,000l.^{4} was destined to pay the marriage portion of the Princess Royal, on her nuptials with the Prince of Orange: the remainder General Oglethorpe^{5} had interest enough in parliament to obtain, for the purpose of carrying over and settling foreign and other Protestants in his new colony of Georgia in America. The project indeed of the trustees for establishing this colony appears to have been equally humane and disinterested: but it is much to be lamented that it should interfere with another of more extensive and lasting utility, which, if it had taken effect by the education of the youth of New England and other colonies, we may venture with great appearance of reason to affirm, would have planted such principles of religion and loyalty among them, as might have gone a good way towards preventing the present unhappy troubles in that part of the world. But to proceed:

3Commons Journal, May 16, 1726.

After having received various excuses, Bishop Gibson, at that time bishop of London (in whose diocese all the West Indies are included) applying to Sir Robert Walpole, then at the head of the treasury, was favoured at length with the following very honest answer: “If you put this question to me,” says Sir Robert, “as a minister, I must and can assure you that the money shall most undoubtedly be paid as soon as suits with public convenience: but if you ask me as a friend, whether Dean Berkeley should continue in America, expecting the payment of 10,100l. I advise him by all means to return home to Europe, and give up his present expectations.” The Dean being informed of this conference by his good friend the Bishop, and thereby fully convinced that the bad policy of one great man had rendered abortive a scheme, whereon he had expended much of his private fortune, and more than seven years of the prime of his life, returned to Europe. Before he left Rhode Island, he distributed what books he had brought with him among the clergy of that province; and immediately after his arrival in London, he returned all the private subscriptions that had been advanced for the support of his undertaking.

In February 1732, he preached before the society for the Propagation of the Gospel in foreign Parts, a sermon since printed at their desire, wherein, from his own knowledge of the state of religion in America, he offers many useful hints towards promoting the noble purposes for which that society was founded.

4Commons Journal, May 10, 1733.

5Ibid.

The same year he gave a more conspicuous proof that he had not mispent the time he had been confined on the other side of the Atlantic, by producing to the world The Minute Philosopher, a masterly performance, wherein he pursues the Free-thinker through the various characters of atheist, libertine, enthusiast, scorner, critic, metaphysician, fatalist, and sceptic; and very happily employs against him several new (O) weapons, drawn from the storehouse of his own ingenious system of philosophy. It is written in a series of dialogues on the model of Plato, a philosopher whom he studied particularly, and whose manner he is thought to have copied with more success than any other that every attempted to imitate him.

We have already related by what means, and upon what occasion, Dr. Berkeley had first the honour of being known to Queen Caroline. This Princess delighted much in attending to philosophical conversations between learned and ingenious men; for which purpose she had, when Princess of Wales, appointed a particular day in the week, when the most eminent for literary abilities at that time in England were invited to attend her Royal Highness in the evening; a practice which she continued after her accession to the throne. Of this company were Doctors Clarke, Hoadley, Berkeley, and Sherlock. Clarke and Berkeley were generally considered as principals in the debates that arose upon those occasions; and Hoadley adhered to the former, as Sherlock did to the latter. Hoadley was no friend to our author: he affected to consider his philosophy and his Bermuda project as the reveries of a visionary. Sherlock (who was afterwards Bishop of London) on the other hand warmly espoused his cause; and particularly when the Minute Philosopher came out, he carried a copy of it to the Queen, and left it to her Majesty to determine whether such a work could be the production of a disordered understanding.

After Dean Berkeley’s return from Rhode Island, the Queen often commanded his attendance to discourse with him on what he had observed worthy of notice in America. His agreeable and instructive conversation engaged that discerning Princess so much in his favour, that the rich Deanery of Down in Ireland falling vacant, he was at her desire named to it, and the King’s letter actually came over for his appointment. But his friend Lord Burlington having neglected to notify the royal intentions in proper time to the Duke of Dorset, then Lord Lieutenant of Ireland, his Excellency was so offended at this disposal of the richest Deanery in Ireland without his concurrence, that it was thought proper not to press the matter any further. Her Majesty upon this declared, that since they would not suffer Dr. Berkeley to be a Dean in Ireland, he should be a Bishop; and accordingly in 1733 the Bishopric of Cloyne becoming vacant, he was by Letters Patent dated March 17, promoted to that see, and was consecrated at St. Paul’s church in Dublin on the 19th of May following by Theophilus Archbishop of Cashel, assisted by the Bishops of Raphoe and Killaloe.

His Lordship repaired immediately to his manse-house at Cloyne, where he constantly resided (except one winter that he attended the business of Parliament in Dublin) and applied himself with vigour to the faithful discharge of all episcopal duties. He revived in his diocese the useful office of Rural Dean which had gone into disuse, visited frequently parochially, and confirmed in the several parts of his see.

He continued his studies however with unabated attention, and about this time engaged in a controversy with the mathematicians of Great Britain and Ireland, which made a good deal of noise in the literary world. The occasion was this: Mr. Addison had given the Bishop an account of their common friend Dr. Garth’s behaviour in his last illness, which was equally unpleasing to both those excellent advocates for revealed religion. For when Mr. Addison went to see the Doctor, and began to discourse with him seriously about preparing for his approaching dissolution, the other made answer, “Surely, Addison, I have good reason not to believe those trifles, since my friend Dr. Halley, who has dealt so much in demonstration, has assured me, that the doctrines of Christianity are incomprehensible, and the religion itself an imposture.” The Bishop therefore took arms against this redoubtable dealer in demonstration, and addressed The Analyst (P) to him, with a view to shewing, that mysteries in faith were unjustly objected to by mathematicians, who admitted much greater mysteries, and even falsehoods in science, of which he endeavoured to prove that the doctrine of fluxions furnished an eminent example. Such an attack upon what had hitherto been looked upon as impregnable, produced a number of warm answers, to which the Bishop replied once or twice.

From this controversy he turned his thoughts to subjects of more apparent utility; and his Queries proposed for the good of Ireland, first printed in 1735, his Discourse addressed to Magistrates, which came out the year following, and his Maxims concerning Patriotism, published in 1750, are equally monuments of his knowledge of mankind, and of his zeal for the service of true religion and his country.

In 1745, during the Scots rebellion, his Lordship addressed a letter to the Roman Catholics of his diocese, and in 1749, another to the Clergy of that persuasion in Ireland under title of A Word to the Wise, written with so much candour and moderation as well as good sense, that those gentlemen, highly to their own honour, in the Dublin Journal of November 18, 1749, thought fit to return “their sincere and hearty thanks to the worthy author, assuring him that they are determined to comply with every particular recommended in his address to the utmost of their power.” They add, that “in every page it contains a proof of the author’s extensive charity; his views are only towards the public good; the means he prescribeth are easily complied with; and his manner of treating persons in their circumstances so very singular, that they plainly shew the good man, the polite gentleman, and the true patriot.” A character this which was so entirely his Lordship’s due, that in the year 1745 that excellent judge of merit, and real friend to Ireland, the late Lord Chesterfield, as soon as he was advanced to government, of his own motion wrote to inform him, that the see of Clogher then vacant, the value of which was double that of Cloyne, was at his service. This offer our Bishop, with many expressions of thankfulness, declined. He had enough already to satisfy all his wishes; and agreeable to the natural warmth of his temper, he had conceived so high an idea of the beauties of Cloyne, that Mr. Pope had once almost determined to make a visit to Ireland on purpose to see a place, which his friend had painted out to him with all the brilliancy of colouring, and which yet to common eyes presents nothing that is very worthy of attention.

The close of a life thus devoted to the good of mankind, was answerable to the beginning of it; the Bishop’s last years being employed in enquiring into the virtues of a medicine, whereof he had himself experienced the good effects in the relief of a nervous cholic, brought on him by his sedentary course of living, and grown to that height, that, in his own words, “it rendered life a burden to him, and the more so, as his pains were exasperated by exercise.” This medicine was no other than the celebrated Tar-water; his thoughts upon which subject he first communicated to the world in the year 1744, in a treatise entitled Siris, a Chain of Philosophical Reflections and Enquiries concerning the Virtues of Tar-water. The author has been heard to declare, that this work cost him more time and pains than any other he had ever been engaged in; a circumstance that will not appear surprising, to such as shall give themselves the trouble of examining into the extent of erudition that is there displayed. It is indeed a chain, which, like that of the poet, resting upon the earth, carries us up to Jupiter, conducting the reader by an almost imperceptible gradation from the phaenomena of tar-water, through the depths of the ancient philosophy, to the sublimest mystery of the Christian religion. It underwent a second impression in 1747, and was followed by Farther Thoughts on Tar-water, published in 1752. This was his last performance for the press, and he survived it but a short time.

In July 1752 he removed, though^{6} in a bad state of health, with his Lady and family to Oxford, in order to superintend the education of one^{7} of his sons, then newly admitted a student at Christchurch. He had taken a fixed resolution to spend the remainder of his days in this city, with a view of indulging the passion for a learned retirement, which had ever strongly possessed his mind, and was one of the motives that led him to form his Bermuda project. But as no body could be more sensible than his Lordship of the impropriety of a Bishop’s non-residence, he previously endeavoured to exchange his high preferment for some canonry or headship at Oxford. Failing of success in this, he actually wrote over to the Secretary of State, to request that he might have permission to resign his Bishoprick, worth at that time at least 1400l. per annum. So uncommon a petition excited his Majesty’s curiosity to enquire who was the extraordinary man that preferred it; being told that it was his old acquaintance Dr. Berkeley, he declared that he should die a Bishop in spite of himself, but gave him full liberty to reside where he pleased.

6He was carried from his landing on the English shore in a horse-litter to Oxford.

7This gentlemen, George Berkeley, second son of the Bishop, proceeded A.M. January 26, 1759, took holy orders, and in August following was presented to the vicarage of Bray in Berkshire. The late Archbishop Secker, who had a high respect for the father’s

The Bishop’s last act before he left Cloyne was to sign a lease of the demesne lands in that neighbourhood, to be renewed yearly at the rent of 200l. which sum he directed to be distributed every year until his return, among poor housekeepers of Cloyne, Youghal, and Aghadda.

At Oxford he lived highly respected by the learned members of that great University, till the hand of Providence unexpectedly deprived them of the pleasure and advantage derived from his residence amongst them. On Sunday evening, January 14, 1753, as he was sitting in the midst of his family, listening to a sermon of Dr. Sherlock’s which his Lady was reading to him, he was seized with what the physicians termed a palsy in the heart, and instantly expired. The accident was so sudden, that his body was quite cold, and his joints stiff, before it was discovered; as the Bishop lay on a couch, and seemed to be asleep, till his daughter, on presenting him with a dish of tea, first perceived his insensibility. His remains were interred at Christchurch, Oxford, where there is an elegant marble monument erected to his memory by his Lady, who is still living, and had during her marriage brought him three sons and one daughter.

As to his person, he was a handsome man, with a countenance full of meaning and benignity, remarkable for great strength of limbs, and, till his sedentary life impaired it, of a very robust constitution. He was however often troubled with the hypochondria, and latterly with that nervous cholic mentioned above.

character, honoured the son with his patronage and friendship, both at the University and afterwards. By his bounty, Dr. Berkeley is now possessed of a canonry of Canterbury, the chancellorship of the collegiate church of Brecknock, and (by exchange for the vicarage of Bray) of the vicarage of Crookham, Berks: to which was added last year, by the Dean and Chapter of Canterbury, the vicarage of East Peckham, Kent. He took the degree of L.L.D. February 12, 1768. In the year 1760, he married the daughter of the Reverend Mr. Frinsham, rector of White-Waltham, Berks, and by this lady hath had issue two sons.

His elder brother, William, died young, in his father’s life-time.

At Cloyne he constantly rose between three and four o’clock in the morning, and summoned his family to a lesson on the base-viol from an Italian master he kept in the house for the instruction of his children; though the Bishop himself had no ear for music. He spent the rest of the morning, and often a great part of the day, in study: his favourite author, from whom many of his notions were borrowed, was Plato. He had a large and valuable collection of books and pictures, which are now the property of his son, the Reverend George Berkeley, L.L.D.

The excellence of his moral character; if it were not so conspicuous in his writings, might be learned from the blessings with which his memory is followed by the numerous poor (Q) of his neighbourhood, as well as from the testimony of his yet surviving acquaintance, who cannot to this day speak of him without a degree of enthusiasm, that removes the air of hyperbole from the well-known line of his friend Mr. Pope:

To Berkeley every Virtue under Heaven.

The inscription on his monument was drawn up by Dr. Markham the present Bishop of Chester, then head master of Westminster School, and is in these terms:

Gravissimo praesuli, Georgio, Episcopo Clonensi: Viro, Seu ingenii & eruditionis laudem, Seu probitatis & beneficentiae spectemus, Inter primos omnium aetatum numerando. Si Christianus fueris, Si amans patriae, Utroque nomine gloriari potes BERKLEIUM vixisse. Obiit annum agens septuagesimum tertium^{8} Natus Anno Christi M.DC.LXXIX. Anna Conjux L.M.P.

8Mistake.

NOTES.

(A) Arithmetica absque Algebra, &c.] This little piece is so far curious, as it shews his early and strong passion for the mathematics, his admiration of those great names in philosophy, Locke and Newton, some of whose positions he afterwards ventured to call in question, and the commencement of his application to those more subtile metaphysical studies, to which his genius was peculiarly adapted.

(B) Theory of Vision.] Of all our author’s works this seems to do the greatest honour to his sagacity; being, as Dr. Reid observes^{9}, the first attempt that ever was made to distinguish the immediate and natural objects of sight, from the conclusions we have been accustomed from infancy to draw from them: a distinction, that gentleman justly adds, from which the nature of vision hath received great light, and by which many phaenomena in optics, before looked upon as unaccountable, have been clearly and distinctly resolved. The boundary is here traced out between the ideas of sight and touch, and it is shewn demonstrably, that though habit hath so connected these two classes of ideas in our minds, that they are called by the same names, and are not without a strong effect of attention to be separated from each other, yet originally they have no such connection; insomuch that a person born blind, and suddenly made to see, would at first be utterly unable to tell how any object that affected his sight would affect his touch, and particularly would not from sight receive any idea of distance, outness, or external space, but would imagine that all the objects he saw were in his eye, or rather in his mind. This last very curious assertion was afterwards, in the year 1728, surprizingly confirmed in the case of a young man born blind, and couched at fourteen years of age by Mr. Cheselden, F. R. S. and Surgeon to her Majesty, whose narrative is so strong in favour of our author’s theory, that we could not forbear presenting the reader with an extract of it from the Philosophical Transactions, N^{o} 402.

‘When he first saw, he was so far from making any judgment about distances, that he thought all objects whatever touched his eyes (as he expressed it) as what he felt touched his skin; and thought no objects so agreeable as those which were smooth and regular, though he could form no judgment of their shape, or guess what it was in any object that was pleasing to him. He knew not the shape of any thing, nor any one thing from another, however different in shape or magnitude; but upon being told what things were whose form he knew before from feeling, he would carefully observe, that he might know them again: but having too many objects to learn at once, he forgot many of them, and as he said, at first he learned to know, and again forgot a thousand things in a day. One particular only, though it may appear trifling, I will relate. Having often forgot which was the cat and which was the dog, he was ashamed to ask, but catching the cat, which he knew by feeling, he was observed to look at her stedfastly, and then setting her down said , So, puss! I shall know you another time.—We thought he soon knew what Pictures represented which were shewed to him, but we found afterwards we were mistaken; for about two months after he was couched, he discovered at once they represented solid bodies, when to that time he considered them only as party-coloured planes, or surfaces diversified with variey of paint: but even then he was no less surprized, expecting the pictures would feel like the things they represented; and was amazed when he found those parts which by their light and shadow appeared now round and uneven, felt only flat like the rest; and asked which was the lying sense, feeling or seeing.—Being shewn his father’s picture in a locket at his mother’s watch, and told what it was, he acknowledged a likeness, but was vastly surprized, asking how it could be that a large face should be expressed in so little room? saying, it should have semmed as impossible to him as to put a bushel of any thing into a pint.’

9Inquiry into the Mind, c. 6. §. 11.

(C) Principles of human Knowledge.] The Theory of Vision is dedicated to Sir John Percival, afterwards Earl of Egmont; the Principles of human Knowledge to the Earl of Pembroke, at that time Lord Lieutenant of Ireland. In the Introduction to this last, the author quarrels with Mr. Locke’s account of abstract ideas and general names. Words, says Mr. Locke, become general, by being made the signs of general ideas, and ideas become general by abstraction, that is, by a power of the mind to conceive separately the several co-existing qualities of an object, and leaving out of the complex idea whatever is peculiar to the individual, and retaining only what is common to it with others of the same kind, to frame an abstract idea, wherein all the particulars of that kind equally partake. According to Berkeley, this operation is impossible. A word becomes general by being made the sign, not of an abstract general idea, but of several particular ideas, any one of which it indifferently suggests to the mind. For example; when we prove any thing concerning motion and extension in general, we do not mean to say that we have an idea of motion without a body moved, or a particular velocity, or of extension without colour, magnitude, &c. but it is implied, that whatever particular motion or extension we consider, the proposition holds equally true concerning it. Again; an idea, which considered in itself is particular, becomes general by being made to represent, or stand for, all other particular ideas of the same sort. The doctrine of abstraction, he says, arose from a mistake concerning the nature of language. It is supposed that the use of language is only to communicate ideas, and that every significant name stands for an idea: since therefore names do not always mark particular ideas it is straightway concluded that they stand for abstract notions. But first, the only end of language is not the communicating of ideas, but the exciting of some passion, &c. which is often accomplished without the intervention of words: and secondly, it is not necessary that names, every time they are used, should excite the ideas they stand for; in reading and discourse they are for the most part used as letters in algebra, wherein although a particular quantity is marked by each letter, it is not requisite that in every step the letter should suggest to our thoughts the quantity it stands for. He infers therefore that we have no ideas but particular ones; and that names do not always, at the time they are used, stand for ideas.

This whole dispute seems to turn upon a single point; whether it is possible for the mind, or pure intellect, to frame to itself ideas which cannot be the object of the imagination, cannot be pictured by fancy; or, in the language of the Peripatetic school, whether there is any ground for the distinction between νoήμαια and ϕαντάσμαια, notions and phantasms. The question must, we think, be answered in the affirmative against Berkeley, by any one who shall attentively read over Cudworth’s Essay on eternal and immutable Morality, IV. 1. §. 8, &c., and IV. 3. §. 6. See also Bolingbroke’s Philosophical Works, Essay I. p. 117, London 1754.

The object of the Principles of human Knowledge, as well as of the Dialogues between Hylas and Philonous, is to prove, that the commonly received notion of the existence of matter is false, and inconsistent with itself; that those things which are called sensible material objects are not external to the mind, but exist in it, and are nothing more than impression made upon our minds by the immediate act of God, according to certain rules, termed laws of nature, from which in the ordinary course of his government he never deviates; that the steady adherence of the supreme Spirit to these rules is what constitutes the reality of things to his creatures, and so effectually distinguishes the ideas perceived by sense from such as are the work of the mind itself or of dreams, that there is no more danger of confounding them together on this hypothesis than on the common supposition of matter. The not attending to this distinction, which however is incalculated over and over by our author, has led many to suppose that Berkeley was an arrant sceptic that rejected the testimony of his senses; when in truth the dispute is not about the reality of our sensations (and so far only the testimony of mere sense extends)—for of this he was as firmly convinced as any body could be, and of the necessity of acting accordingly—but concerning the causes of those sensations, whether they proceed from a set of insensible material beings without us, or immediately from the Creator himself. His principal argument against the existence of those material beings may be reduced to this syllogism:

Whatever is immediately perceived by sense, is an idea.

Sensible things are things immediately perceived by sense.

Therefore sensible things are ideas; and consequently exist only in the mind.

For the proof of the second proposition he appeals to the feelings of his reader, and asks, whether by what he calls sensible things he means any thing else but the things he immediately perceives by sense? whether, for instance, when he says this table exists, he means any thing more than that he sees and feels it, that is, perceives it by his senses? If you answer, that you are indeed immediately conscious only of the sensation, but that reason obliges you to infer from thence the existence of an external being which is the cause of it, he joins issue with you and says, the question then is only whether this external cause is active or inactive, spirit or matter. Now it is a contradiction to suppose that unthinking matter can be the cause of ideas: for causality supposes action: action must consist either in motion or volition: it cannot consist in the former, because motion is a sensible quality, i. e. an idea; but all our ideas are passive, inert, including nothing of action in them: therefore nothing can be active, and consequently nothing can be a cause without volition. See Hylas, page 265, London 1734.—But, not to mention that a thing is here asserted which no force of argument will ever make out, that a quality of a body is the same thing with an idea in the mind, it is obvious to observe, that the whole is merely an argument ab ignorantiâ, ground on our inability of shewing the manner by which body operates on spirit; an inability, however, which cannot take away that of whose existence we have otherwise good reason to be satisfied. This good reason then appears to us to be simply the very same that Dr. Reid first pointed out, the constitution of our nature (we should call it instinct, if such an out cry had not lately been raised against that word) which antecedent to, and independent of all reasoning about the matter, compels us to believe the existence of a number of beings without us both animate and inanimate, with as strong and invincible a faith as we believe the existence of ourselves and our own sensations. If we are not to trust this constitution in the former case, no reason can be given why we should take its word for the truth of our own consciousness. “It is folly,” saith Mr. Locke, “to expect demonstration in every thing.” Descartes attempted to prove his own existence. Berkeley could find no proof at all for the existence of matter, and but a probability (though he allows it to be a high probability, see Hylas, p. 297, l. 1.) for that of his fellow creatures. The success of these enquiries may serve to verify Mr. Locke’s observation.

When the Principles of human Knowledge were first published, the ingenious author sent copies of the work to Dr. Clarke and Mr. Whiston. What effect it produced upon the latter, the reader may possibly be entertained with learning from his own words: Memoirs of Dr. Clarke, page 79—

‘And perhaps it will not be here improper, by way of caution, to take notice of the pernicious consequence such metaphysical subtilties have sometimes had, even against common sense and common experience, as in the cases of those three famous men, Mons. Leibnitz, Mr. Locke, and Mr. Berkeley.—[The first in his pre-established Harmony: the second in the dispute with Limborch about human Liberty.]—And as to the third named, Mr. Berkeley, he published A. D. 1710, at Dublin, this metaphysick notion, that matter was not a real thing; nay, that the common opinion of its reality was groundless, if not ridiculous. He was pleased to send Dr. Clarke and myself, each of us, a book. After we had both perused it I went to Dr. Clarke, and discoursed with him about it to this effect; that I, being not a metaphysician, was not able to answer Mr. Berkeley’s subtile premises, though I did not at all believe his absurd conclusion. I therefore desired that he, who was deep in such subtilties, but did not appear to believe Mr. Berkeley’s conclusions, would answer him; which talk he declined. I speak not these things with intention to reproach either Mr. Locke or Dean Berkeley.—I own the latter’s great abilities in other parts of learning; and to his noble design of settling a College in or near the West Indies, for the instruction of the natives in civil arts and in the principles of Christianity, I heartily wish all possible success. It is the pretended metaphysick science itself, derived from the sceptical disputes of the Greek philosophers, not those particular great men who have been unhappily imposed on by it, that I complain of. Accordingly when the famous Milton had a mind to represent the vain reasonings of wicked spirits in Hades, he describes it by their endless train of metaphysicks thus:

‘Others apart sat on a hill retired,’ &c.

Par. Lost, II. 557–561.

(D) Lord Berkeley of Stratton.] Swift presented him to this nobleman as his relation, with this humorous speech, that he was good for something. Observations on Orrery’s Life of Swift, page 122, 8vo. 1734. In a letter to Stella, April 12, 1713, the Dean speaks thus: ‘I went to court to-day on purpose to present Mr. Berkeley, one of your Fellows of Dublin College, to Lord Berkeley of Stratton. That Mr. Berkeley is a very ingenious man, and a great philosopher; and I have mentioned him to all the ministers, and I have given them some of his writings, and I will favour him as much as I can. This I think I am bound to in honour and conscience, to use all my little credit towards helping forward men of worth in the world.’

(E) In August 1714.] Towards the close of this year he had a fever, in describing the event of which to his friend Swift, Dr. Arbuthnott cannot forbear indulging a little of that pleasantry on Berkeley’s system, with which it has frequently since been treated by such as could not, or would not, be at the pains to acquire a thorough knowledge of it. ‘October 19, 1714, poor philosopher Berkeley has now the idea of health, which was very hard to produce in him; for he had an idea of a strange fever on him so strong, that it was very hard to destroy it by introducing a contrary one.’

(F) To recollect his observations on Sicily.] One discovery of his only in that country his friends remember to have heard from him; that in the town by Virgil called palmosa Selinus, Æneid III. 705, he found the ruins of a most magnificent temple dedicated to Pollux, to which the natives have given the appellation of Tempio di Pulci, corruptly for Polluci; and the territory adjoining has from thence taken the name of Terra de li Pulci.

(G) His Letters to Pope and Arbuthnott.] The first is in Pope’s Works, Vol. VIII. p. 259, London 1757. The other is in the Philosophical Transactions, N^{o} 354. We shall make no apology for giving them both a place in this narrative.

‘Leghorn, May 1, 1714.

‘As I take ingratitude to be a greater crime than impertinence, I chuse rather to run the risque of being thought guilty of the latter, than not to return you my thanks for a very agreeable entertainment you just now gave me. I have accidentally met with your Rape of the Lock here, having never seen it before. Style, painting, judgment, spirit, I had already admired in other of your writings; but in this I am charmed with the magic of your invention, with all those images, allusions, and inexplicable beauties, which you raise so surprisingly, and at the same time so naturally, out of a trifle. And yet I cannot say that I was more pleased with the reading of it, that I am with the pretext it gives me to renew in your thoughts the remembrance of one who values no happiness beyond the friendship of men of wit, learning, and good-nature.

‘I remember to have heard you mention some half-formed design of coming to Italy. What might we not expect from a muse that sings so well in the bleak climate of England, if she felt the same warm fun, and breathed the same air with Virgil and Horace!

‘There are here an incredible number of Poets that have all the inclination, but want the genius, or perhaps the art of the ancients. Some among them who understand English, begin to relish our authors; and I am informed that at Florence they have translated Milton into Italian verse. If one who knows so well how to write like the old Latin poets came among them, it would probably be a means to retrieve them from their cold trivial conceits, to an imitation of their predecessors.

‘As merchants, antiquaries, men of pleasure, &c. have all different views in travelling, I know not whether it might not be worth a poet’s while to travel, in order to store his mind with strong images of nature.

‘Green fields and groves, flowery meadows and purling streams, are no where in such perfection as in England: but if you would know lightsome days, warm suns, and blue skies, you must come to Italy, and to enable a man to describe rocks and precipices, it is absolutely necessary that he pass the Alps.

‘You will easily perceive that it is self-interest makes me so fond of giving advice to one who has no need of it. If you came into these parts I should fly to see you. I am here (by the favour of my good friend the Dean of St. Patrick’s) in quality of Chaplain to the Earl of Peterborough, who about three months since left the greatest part of his family in this town. God knows how long we shall stay here. I am,

‘Your, &c.’

‘April 17, 1717, with much difficulty I reached the top of Mount Vesuvius, in which I saw a vast aperture full of smoke, which hindered the seeing its depth and figure. I heard within that horrid gulf certain odd sounds, which seemed to proceed from the belly of the mountain; a sort of murmuring, sighing, throbbing, churning, dashing (as it were) of waves, and between whiles a noise like that of thunder or cannon, which was constantly attended with a clattering like that of tiles falling from the tops of houses on the streets. Sometimes as the wind changed the smoke grew thinner, discovering a very ruddy flame, and the jaws of the pan or crater streaked with red and several shades of yellow. After an hour’s stay the smoke, being moved by the wind, gave us short and partial prospects of the great hollow, in the flat bottom of which I could discern two furnaces almost contiguous: that on the left, seeming about three yards in diameter, glowed with red flame, and threw up red-hot stones with a hideous noise, which, as they fell back, caused the fore-mentioned clattering. May 8, in the morning, I ascended to the top of Vesuvius a second time, and found a different face of things. The smoke ascending upright gave a full prospect of the crater, which, as I could judge, is about a mile in circumference, and an hundred yards deep. A conical mount had been formed since my last visit, in the middle of the bottom: this mount, I could see, was made of the stones thrown up and fallen back again into the crater. In this new hill remained the two mounts of furnaces already mentioned: that on our left was in the vertex of the hill which had formed around it, and raged more violently than before, throwing up every three or four minutes with a dreadful bellowing a vast number of red-hot stones, sometimes in appearance above a thousand, and at least 3000 feet higher than my head as I stood upon the brink: but there being little or no wind, they fell back perpendicularly into the crater, increasing the conical hill. The other mouth to the right was lower in the side of the same new formed hill: I could discern it to be filled with red-hot liquid matter, like that in the furnace of a glass-house, which raged and wrought as the waves of the sea, causing a short abrupt noise like what may be imagined to proceed from a sea of quicksilver dashing among uneven rocks. This stuff would sometimes spew over and run down the convex side of the conical hill; and appearing at first red-hot it changed colour, and hardened as it cooled, shewing the first rudiments of an eruption, or if I may so say, an eruption in miniature. Had the wind driven in our faces, we had been in no small danger of stifling by the sulphureous smoke, or being knocked on the head by lumps of molten minerals which we say had sometimes fallen on the brink of the crater, upon those shots from the gulf at bottom. But as the wind was favourable, I had an opportunity to survey this odd scene for above an hour and a half together, during which it was very observable, that all the volleys of smoke, flame, and burning stones, came only out of the hole to our left, while the liquid stuff in the other mouth wrought and overflowed, as hath been already described. June 5, after a horrid noise, the mountain was seen at Naples to spew a little out of the crater. The same continued the 6th. The 7th, nothing was observed till within two hours of night, when it began a hideous bellowing, which continued all that night and the next day till noon, causing the windows, and, as some affirm, the very houses in Naples to shake. From that time it spewed vast quantities of molten stuff to the South, which streamed down the side of the mountain like a great pot boiling over. This evening I returned from a voyage through Apulia, and was surprized, passing by the North side of the mountain, to see a great quantity of ruddy smoke lie along a huge tract of sky over the river of molten stuff, which was itself out of sight. The 9th, Vesuvius raged less violently: that night we saw from Naples a column of fire shoot between whiles out of its summit. The 10th, when we thought all would have been over, the mountain grew very outragious again, roaring and groaning most dreadfully. You cannot form a juster idea of this noise in the most violent fits of it, than by imagining a mixed sound made up of the raging of a tempest, the murmur of a troubled sea, and the roaring of thunder and artillery, confused all together. It was very terrible as we heard it in the further end of Naples, at the distance of above twelve miles: this moved my curiosity to approach the mountain. Three or four of us got into a boat, and were set ashore at Torre del Greco, a town situate at the foot of Vesuvius to the South-west, whence we rode four or five miles before we came to the burning river, which was about midnight. The roaring of the volcano grew exceedingly loud and horrible as we approached. I observed a mixture of colours in the cloud over the crater, green, yellow, red, and blue; there was likewise a ruddy dismal light in the air over that tract of land where the burning river flowed; ashes continually showered on us all the way from the sea-coast: all which circumstances, set off and augmented by the horror and silence of the night, made a scene the most uncommon and astonishing I ever saw, which grew still more extraordinary as we came nearer the stream. Imagine a vast torrent of liquid fire rolling from the top down the side of the mountain, and with irresistible fury bearing down and consuming vines, olives, figtrees, houses; in a word, every thing that stood in its way. This mighty flood divided into different channels, according to the inequalities of the mountain: the largest stream seemed half a mile broad at least, and five miles long. The nature and consistence of these burning torrents hath been described with so much exactness and truth by Borellus in his Latin treatise of Mount Ætna, that I need say nothing of it. I walked so far before my companions up the mountain, along the side of the river of fire, that I was obliged to retire in great haste, the sulphurous steam having surprized me, and almost taken away my breath. During our return, which was about three o’clock in the morning, we constantly heard the murmur and groaning of the mountain, which between whiles would burst out into louder peals, throwing up huge spouts of fire and burning stones, which falling down again, resembled the stars in our rockets. Sometimes I observed two, at others three distinct columns of flame; and sometimes one vast one that seemed to fill the whole crater. These burning columns and the fiery stones seemed to be shot 1000 feet perpendicular above the summit of the volcano. The 11th at night I observed it from a terrass in Naples, to throw up incessantly a vast body of fire, and great stones to a surprizing height. The 12th in the morning it darkened the sun with ashes and smoke, causing a sort of eclipse. Horrid bellowings this and the foregoing day were heard at Naples, whither part of the ashes also reached. At night I observed it throw up flame as on the 11th. On the 13th the wind changing, we saw a pillar of black smoke shot upright to a prodigious height: at night I observed the mount cast up fire as before, though not so distinctly because of the smoke. The 14th, a thick black cloud hid the mountain from Naples. The 15th in the morning the court and walls of our house in Naples were covered with ashes. The 16th, the smoke was driven by a Westerly wind from the town to the opposite side of the mountain. The 17th the smoke appeared much diminished, fat, and greasy. The 18th the whole appearance ended; the mountain remaining perfectly quiet without any visible smoke or flame. A gentleman of my acquaintance, whose window looked toward Vesuvius, assured me that he observed several flashes, as it were of lightening, issue out of the mouth of the volcano. It is not worth while to trouble you with the conjectures I have formed concerning the cause of these phænomena, from which I observed in the Lacus Amsancti, the Solfatara, &c. as well as in Mount Vesuvius. One thing I may venture to say, that I saw the fluid matter rise out of the centre of the bottom of the crater, out of the very middle of the mountain, contrary to what Borellus imagines, whose method of explaining the eruption of a volcano by an inflexed syphon and the rules of hydrostatics, is likewise inconsistent with the torrent’s flowing down from the very vertex of the mountain. I have not seen the crater since the eruption, but design to visit it again before I leave Naples. I doubt there is nothing in this worth shewing to the Society; as to that, you will use your discretion.

‘E. (it should be G.) BERKELEY.’

(H) A Tract De Motu.] The principal positions in this treatise are,

1. That neither the beginning nor the communication of motion can justly be ascribed to body, which is wholly incapable of action, but must be referred to spirit only, and ultimately to the supreme Spirit, the fountain of all things. 2. That pure space is a mere figment of philosophers; space not being absolute, but relative to the bodies comprehended in it: so that if these were annihilated, space would perish along with them, like all other relations, which connot be conceived to exist without their correlatives.

(I) A Proposal for Converting the Savage Americans.] With this proposal he carried a letter of recommendation from Dean Swift to Lord Carteret, Lieutenant of Ireland, which deserves a place here, both because it contains a number of particulars of our Author’s life, and is besides a proof, as well of the friendly temper of the writer, as of his politeness and address:

“September 3, 1724.—There is a gentleman of this kingdom just gone for England; it is Doctor George Berkeley, Dean of Derry, the best preferment among us, being worth about 1100l. a year. He takes the Bath in his way to London, and will of course attend your Excellency, and be presented I suppose by his friend my Lord Burlington; and, because I believe you will chuse out some very idle minutes to read this letter, perhaps you may not be ill entertained with some account of the man and his errand. He was a Fellow in the University here, and going to England very young, about thirteen years ago, he became the founder of a sect there called the Immaterialists, by the force of a very curious book on that subject: Dr. Smalridge and many other eminent persons were his proselytes. I sent him Secretary and Chaplain to Sicily with my Lord Peterborough; and upon his Lordship’s return, Dr. Berkeley spent above seven years in travelling over most parts of Europe, but chiefly through every corner of Italy, Sicily, and other islands. When he came back to England he found so many friends, that he was effectually recommended to the Duke of Grafton, by whom he was lately made Dean of Derry. Your Excellency will be frighted when I tell you all this is but an introduction; for I am now to mention his errand. He is an absolute philsopher with regard to money, titles and power; and for three years past has been struck with a notion of founding a University at Bermuda, by a charter from the crown. He hath seduced several of the hopefullest young clergymen and others here, many of them well provided for, and all of them in the fairest way of preferment: but in England his conquests are greater, and I doubt will spread very far this winter. He shewed me a little tract which he designs to publish, and there your Excellency will see his whole scheme of a life academico-philosophical (I shall make you remember what you were) of a college founded for Indian scholars and missionaries, where he most exorbitantly proposeth a whole hundred pounds a year for himself, forty pounds for a fellow, and ten for a student. His heart will break if his Deanery be not taken from him, and left to your Excellency’s disposal. I discourage him by the coldness of courts and ministers, who will interpret all this as impossible, and a vision; but nothing will do. And therefore I humbly entreat your Excellency either to use such persuasions as will keep one of the first men in this kingdom for learning and virtue quite at home, or assist him by your credit to compass his romantic design; which however is very noble and generous, and directly proper for a great person of your excellent education to encourage.”

(K) Certain Lands in St. Christopher’s.] “The island of St. Christopher’s,” saith Anderson, History of Commerce, Vol. II. “having been settled on the very same day and year by both England and France, A. D. 1625, was divided equally between the two nations. The English were twice driven out from thence by the French, and as often re-possessed themselves of it. But at length, in the year 1702, General Coddrington, Governor of the Leeward Islands, upon advice received that war was declared by England against France, attacked the French part of the island, and mastered it with very little trouble. Ever since which time that fine island has been solely possessed by Great Britain, having been formally conceded to us by the treaty of Utrecht.” The lands, therefore, which had belonged to the French planters, by this cession became the property of his Britannic Majesty. The first proposals for purchasing these lands were made to the Lords of Trade in 1717: see Journal of the British Commons. After which, the affair seems to have been forgotten, till it was mentioned by Berkeley to Sir Robert Walpole in 1726.

(L) To King George I.] It was the custom of this prince to unbend his mind in the evening by collecting together a company of philosophical foreigners, who discoursed in an easy and familiar manner with each other, entirely unrestrained by the presence of his Majesty, who generally walked about, or sat in a retired part of the chamber. One of this select company was Altieri; and this gave him an opportunity of laying his friend’s proposal before the King.

(M) A Copy of Verses.] They run thus:

The Muse, disgusted at an age and clime Barren of every glorious theme, In distant lands now waits a better time Producing subjects worthy fame;

In happy climes, where from the genial sun And virgin earth such scenes ensue, The force of art by nature seems outdone, And fancied beauties by the true:

In happy climes, the feat of innocence, Where nature guides, and virtue rules; Where men shall not impose for truth and sense The pedantry of courts and schools.

There shall be sung another golden age,

The rise of empire and arts,

The good and great inspiring epic rage, The wisest heads and noblest hearts.

Not such as Europe breeds in her decay; Such as she bred when fresh and young, When heavenly flame did animate her clay, By future poets shall be sung.

Westward the course of empire takes its way:

The four first acts already past,

A fifth shall close the drama with the day:

Time’s noblest offspring is the last.

(N) His intended Library.] Dr. Berkeley has left a monument of his skill in architecture in some very elegant designs of this and other parts of his new college, drawn by himself, which were formerly in the possession of the Rev. Dr. Raymond, Vicar of Trim, and are now the property of his granddaughter Mrs. Ewing, widow of Mr. Thomas Ewing, an ingenious and spirited Bookseller of Dublin lately deceased, whose death has deprived the public of what has long been much wanted, a complete edition of Bishop Berkeley’s works in 4to.

(O) Several new Arguments from his own System.] The most remarkable of these is a proof of the existence of God, from the principles established in the New Theory of Vision. It is an argument ad hominem, in answer to a somewhat unreasonable demand of a minute philosopher, who insists upon having the same evidence for the being of a God, that he has for that of his fellow mortals, to wit, the use of speech. It is the arbitrary use, saith the objector, of sensible signs, which have no similtude or necessary connection with the things signified by them, and which yet suggest to my mind an endless variety of things serving to inform me, and direct my conduct both present and future; it is this principally that convinces me of the existence of another intelligent, thinking person, speaking to me: shew me that God speaks to man in the same manner by the intervention of arbitrary signs, no matter which sense they enter into the mind, whether by the ear of the eye, provided they have no necessary connection with the things they suggest— such a proof as this only will content me. Such a proof as this you actually have, replies his opponent, in the connection you find between the objects of your sight and touch, which is perfectly arbitrary, and as much learned from experience as the signification of words is. A man born blind would not, at first receiving his sight, be able from the visible appearance or figure of an object to form any judgment of its tangible distance or figure, no more than a Chinese upon first hearing the words man and tree would think of the things signified by them: in both cases there must be time and experience to acquire a habit of knowing the connection between the sign and the thing signified, that is, of understanding the language whether of the eyes or ears. If the connection appears necessary between the objects of sight and touch, and not so between words and things, the reason is, because verbal languages are particular; that of vision is universal, learned from infancy, and practised through life: whence men are more apt to confound the sign with the thing signified in one kind of language than in the other. If the use then of sensible, arbitrary signs, having no necessary coherence with the things suggested by them, is allowed to be a proof of the being and intelligence of him who employs them, you have as much reason to think the universal agent, or God, speaks to your eyes, as you can have for thinking any particular person speaks to your ears.—The compliment Cicero pays to an argument of Plato’s in favour of the soul’s immortality, may with more propriety perhaps be applied to this new and subtle proof of the being of a God: Licet concurrant plebeii omnes philosophi (sic enim ii qui a Platone, & Socrate, & ab eâ familiâ dissident, appellandi videntur) non modo nihil umquam tam eleganter explicabunt, sed ne hoc quidem ipsum quàm subtiliter conclusum sit, intelligent: Though all the vulgar philosophers of the world were to join their powers together, so far from being able to imitate the elegance of this reasoning, they will never have sense enough to comprehend the ingenuity of it.

(P) The Analyst.] The Bishop’s chief objections to the doctrine of fluxions may be comprized under these two heads:

I. That the object (viz. fluxions) was inconceivable.

II. That the arguments, brought to prove the truth of the fundamental proposition were fallacious and inconclusive.

I. It had been said, that fluxions are the velocities with which quantities vary or are generated; and that they are as the momenta, or in the first proportion of the nascent increments. He objects, that he cannot form any clear idea of these momenta or nascent increments, and still less so of the abstracted velocities of such nascent, imperfect entities: and elsewhere he insists upon the impossibility of conceiving velocity without motion, or motion without time and space.

Again; a second fluxion, being called a fluxion of a fluxion, he calls it the velocity of a velocity, the incipient celerity of an incipient celerity, a nascent augment of a nascent augment; and as such, he pronounces the clear conception of it to be impossible. Now if a second fluxion be inconceivable, what are we to think of third, fourth, fifth fluxions, and so onward without end?

II. A principal proposition in the doctrine of fluxions is the rule for finding the momentum or fluxion of a product or power. Of this rule he found two demonstrations in Newton’s works: the first is in the 2d lemma of the second book of the Principles, concerning which he observes, that such reasoning as this for demonstration, nothing but the obscurity of the subject could have encouraged the great author of the fluxionary method to put upon his followers, and nothing but an implicit deference to authority could move them to admit. The other demonstration of this rule is found in the introduction to Newton’s Quadratures; to which the Bishop objects in the same peremptory manner, insisting that it is fallacious, because it proceeds on two contrary hypotheses: for he says, when the second of these hypotheses is made, the first is destroyed, and together with it, all consequences derived from it; whereas the conclusion is drawn from the second hypothesis, and a consequence of the first, jointly.

Supposing these objections to be valid, he brings some heavy charges against the illustrious author of the doctrine: That he placed his fluxions in various lights, and shifted his notions of them; that terms, which should be steadily used, were ambiguous; that he employed various arts and devices, &c.

Several answers to the Analyst soon appeared. Of these, if we except Colson’s commentary then first subjoined to an edition of Newton’s fluxions, the principal were set forth by a gentleman who concealed himself under the name of Philalethes Cantabrigiensis, but who is generally supposed to be Dr. Jurin, and by Benjamin Robins, Esq. The compass of this note will not permit us to do justice to the answers of these two gentlemen: it is sufficient to observe, that though they differed from each other about the meaning of some passages in Newton’s writings, yet they vindicated, to the entire satisfaction of the mathematical reader, the doctrine of fluxions from all the objections that had been advanced against it by the Bishop. Philalethes very soon published a Letter to the author of the Analyst, under the title of Geometry no Friend to Infidelity, in which, after some sharp animadversions on his Lordship, he endeavours to point out to him his mistakes, and answers several of his principal objections. The following year (1735) the Bishop, in a reply intitled A Defence of Free-Thinking in Mathematics, declared himself still of the same opinion, the arguments of his antagonist being futile, and several objections remaining unanswered, and consequently in full force. This drew from Philalethes a second answer in June 1735, styled The minute Mathematician: or the Free-thinker no just Thinker, wherein the defects of the first paper were supplied, the nature of fluxions of all orders explained in the clearest manner, and the former arguments repeated, illustrated, and cleared from the objections the Bishop had raised against them. And here this controversy ended.

Mr. Robins, in the year 1735, published his answer, intitled A Discourse concerning the Nature and Certainty of Sir Isaac Newton’s Methods of Fluxions, and of prime and ultimate Ratios. He proceeded on a different plan from Philalethes; for, without taking any notice of the Analyst or his objections, he delivered the principles of the method of fluxions in such a manner, as that it should not even in appearance be liable to these or any other objections, his demonstrations not yielding in accuracy to those of the antient geometricians, so much celebrated on that very account.

But whatever matter of reprehension mathematicians may find in this hostile attempt of our Bishop, it must be acknowledged they have reaped no inconsiderable advantage from it, inasmuch as it gave rise to the most complete Treatise of Fluxions^{10} that hath been yet presented to the public, in which the whole doctrine is delivered with more precision and fullness than ever was done before, or probably than ever would be done, if no attack had been made upon it.

(Q) By the Poor of his Neighbourhood.] One instance of his attention to his poor neighbours may deserve relating. Cloyne is an Irish village; it is not reasonable therefore to expect much industry or ingenuity in the inhabitants. Yet whatever article of cloathing they could possibly manufacture there, the Bishop would have from no other place; and chose to wear ill cloaths, and worse wigs, rather than suffer the poor of the town to remain unemployed.

FINIS.

10McLaurin’s.

POSTSCRIPT.

The author is this memoir was too far removed from the press to correct in time a mistake, which the reader it is hoped will the more readily pardon, as it submits to his perusal two entertaining letters of Bishop Berkeley instead of one. In note (G) a letter dated Leghorn, May 1, 1714, is inserted from Pope’s Works, instead of the following, which bears date,

‘Naples, Oct 22, 1717.

‘I have long had it in my thoughts to trouble you with a letter, but was discouraged for want of something that I could think worth sending fifteen hundred miles. Italy is such an exhausted subject, that I dare say you would easily forgive my saying nothing of it; and the imagination of a Poet is a thing so nice and delicate, that it is no easy matter to find out images capable of giving pleasure to one of the few, who (in any age) have come up to that character. I am nevertheless lately returned from an island, where I passed three or four months; which, were it set out in its true colours, might, methinks, amuse you agreeably enough for a minute or two. The island Inarime is an epitome of the whole earth, containing within the compass of eighteen miles a wonderful variety of hills, vales, ragged rocks, fruitful plains, and barren mountains, all thrown together in a most romantic confusion. The air is in the hottest season constantly refreshed by cool breezes from the sea. The values produce excellent wheat and Indian corn, but are mostly covered with vineyards, intermixed with fruit-trees; besides the common kinds, as cherries, apricots, peaches, &c. they produce oranges, limes, almonds, pomegranates, figs, water-melons, and many other fruits unknown to our climates, which lie every where open to the passenger. The hills are the greater part covered to the top with vines, some with chesnut groves, and others with thickets of myrtle and lentiscus. The fields on the northern side are divided by hedge-rows of myrtle. Several fountains and rivulets add to the beauty of this landscape, which is likewise set off by the variety of some barren spots, and naked rocks. But that which crowns the scene is a large mountain, rising out of the middle of the Islands (once a terrible Volcano, by the ancients called Mons Epomeus); its lower parts are adorned with vines and other fruits: the middle affords pasture to flocks of goats and sheep, and the top is a sandy pointed rock, from which you have the finest prospect in the world, surveying at one view, besides several pleasant islands lying at your feet, a tract of Italy about three hundred miles in length, from the promontory of Antium to the cape of Palinurus: the greater part of which hath been sung by Homer and Virgil, as making a considerable part of the travels and adventures of their two Heroes. The islands Caprea, Prochyta, and Parthenope, together with Cajeta, Cumae, Monte Miseno, the habitations of Circe, the Syrens, and the Læstrigones, the bay of Naples, the promontory of Minerva, and the whole Campagnia felice, make but a part of this noble landscape; which would demand an imagination as warm, and numbers as flowing as your own, to describe it. The inhabitants of this delicious isle, as they are without riches and honours, so are they without the vices and follies that attend them; and were they but as much strangers to revenge, as they are to avarice and ambition, they might in fact answer the poetical notions of the golden age. But they have got, as an alloy to their happiness, an ill habit of murdering one another on slight offences. We had an instance of this the second night after our arrival, a youth of eighteen being shot dead by our door: and yet by the sole secret of minding our own business, we found a means of living securely among these dangerous people. Would you know how we pass the time at Naples? Our chief entertainment is the devotion of our neighbours: besides the gaiety of their churches (where folks go to see what they call una bella devotione, i. e. a sort of religious opera) they make fireworks almost every week, out of devotion; the streets are often hung with arras, out of devotion; and (what is still more strange) the ladies invite gentlemen to their houses, and treat them with music and sweetmeets, out of devotion; in a word, were it not for this devotion of its inhabitants, Naples would have little else to recommend it beside the air and situation. Learning is in no very thriving state here, as indeed no where else in Italy; however, among many pretenders, some men of taste are to be met with. A friend of mine told me not long since, that being to visit Salvini at Florence, he found him reading your Homer; he liked the notes extremely, and could find no other fault with the version, but that he thought it approached too near a paraphrase; which shews him not to be sufficiently acquainted with our language. I wish you health to go on with that noble work, and when you have that, I need not wish you success. You will do me the justice to believe, that whatever relates to your welfare is sincerely wished by

‘Your, &c.’